real or imaginary solutions to each equation
I put prens. around all that is squared, I hope this helps, I did not know anyother way to write it.
94. (sqrt 7x + 29) = x+3

Answer by tutorcecilia(2152)

(sqrt(7x + 29))^2 = (x+3)^2[square both sides]
7x+29=x^2+6x+9 [set the equation equal to zero]
0=x^2+6x-7x+9-29
0= x^2-x-20 [Factor]
0=(x-5)(x+4)[set each factor equal to zero and solve]
x-5=0
x=5
or
x+4=0
x=-4
Solution: x=5 or x=-4